OpenStudy (anonymous):

WILL FAN AND MEDAL Select the polynomial that is a perfect square trinomial. 49x2 − 8x + 16 4a2 − 10a + 25 25b2 − 5b + 10 16x2 − 8x + 1

OpenStudy (anonymous):

a number after the variable is an exponent

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

@Maxmius @mbma526

OpenStudy (taylor0402):

here is a reference to help you file:///C:/Users/Owner/Downloads/StudyGuide.IA.pdf

OpenStudy (anonymous):

I can't click the link @Taylor0402

OpenStudy (taylor0402):

copy and paste if u doof

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

allo ospreytriple :)

OpenStudy (anonymous):

Hi. You need some help? Firstly, for a trinomial to be a perfect square, both the first and last terms must be perfect squares. For example\[\left( 2x+3 \right)\left( 2x+3 \right) = 4x^2+12x+9\]Both the first term, 4x^2, and the last term, 9, are perfect squares. Understand so far?

OpenStudy (anonymous):

yes @ospreytriple

OpenStudy (anonymous):

This will eliminate one of the answers.

OpenStudy (anonymous):

Now, to pin it down, The coefficient of the middle term (the coefficient of the 'x' term) must be equal to twice the product of the square roots of the coefficient of the first term and the third term. In my example\[4x^2+12x+9\]the square root of the coefficient of the first term is 2 and the square root of the third term is 3. Twice their product is 2(2 x 3) = 12. OK?

OpenStudy (anonymous):

And to make it even trickier, you have to consider that the negative square root may also be a possibility.

OpenStudy (anonymous):

Take the first answer given for example. The square root of 49 is +7 or -7 and the square root of 16 is +4 or -4. So, if it is a perfect square the coefficient of the middel term must be one of 2(-7 x 4) 2(-7 x -4) 2(7 x 4) 2(7 x -4) If it is not one of these, the it is not a perfect square trinomial.

OpenStudy (anonymous):

BTW these are the answers 49x2 − 8x + 16 4a2 − 10a + 25 25b2 − 5b + 10 16x2 − 8x + 1

OpenStudy (anonymous):

@ospreytriple is it B?

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

Yes. Well done.

OpenStudy (anonymous):

THX SO MUCH

OpenStudy (anonymous):

Oops. sorry, it's not B.

OpenStudy (anonymous):

The middle coefficient must be TWICE the product of the two square roots.

OpenStudy (anonymous):

So d

OpenStudy (anonymous):

So, for the possibilities for the coefficient of the middle term are 2(-2 x 5) 2(-2 x -5) 2(2 x 5) 2(2 x -5)

OpenStudy (anonymous):

Yes, D is correct. Sorry for the mixup.

OpenStudy (anonymous):

its ok